# Perils of Label Indeterminacy: A Case Study on Prediction of Neurological Recovery After Cardiac Arrest

### v1

@@ 4.5 Results: Predictions on Label Indeterminacy Set
Table 3: Number of patients with predicted likelihood to recover of $\leq$1in one model (rows) and $>$10% in another (columns). This shows that the ordering of predictions changes between models.
@@
Figure 5: Predictions for two WLST patients with strong disagreement between models. This disagreement is entirely due to unverifiable design choices made in estimating and incorporating labels for WLST patients.
5 Implications
@@
Deploying AI systems in high-stakes domains introduces specific ethical challenges that stem from label indeterminacy. A key concern is the arbitrariness of decisions these systems may inform, wherein different and unjustifiable methodological choices for handling indeterminate labels can yield outcomes lacking any principled grounding [12, 5]. For instance, a system might predict no chance of recovery for a patient under one label-handling method, while another method suggests a positive recovery potential for that same patient. In a healthcare context, this arbitrariness may influence whether life-sustaining therapies are withdrawn, potentially resulting in premature life-or-death decisions. This echoes concerns stemming from predictive multiplicity [46, 13]. While there is a similarity insofar as arbitrary choices yield different predictions, there is a differentiating factor: in the case of label indeterminacy, the issue does not stem from underspecification, under which one problem formulation with a known true label accepts multiple solutions [13]. Instead, it stems from an ill-posed problem formulation, which relies on unknowable information, and in the face of it makes an arbitrary choice as to what that information may be. The “ground truth” is chosen to be one among many options, without knowledge of which one is more likely to be correct, but with full knowledge that at most one of them may be correct. As a result, the differences in prediction stem from some models relying on incorrect information. This poses unique ethical challenges for the design and deployment of these systems.
Crucially, in current practice there is often an implicit and unchallenged assumption that one method of estimating labels is correct, with alternative approaches rarely compared. This narrow focus obfuscates the indeterminate nature of the problem. Moreover, the opacity of assumptions implicit in this selection process can obscure mechanisms behind AI recommendations. For patients and providers, this is particularly problematic as it veils an important dimension of uncertainty of the AI recommendations. This can, among other risks, foster an over-reliance on these recommendations [8, 64]. Without a clear understanding of the uncertainties and limitations, patients and providers might trust the AI recommendations uncritically, potentially introducing new and consequential risks in care delivery.
@@
This also carries implications for research and practice on human-AI collaboration [25, 3, 62]. Recent work in this field has advocated for a focus on appropriate reliance when evaluating AI-assisted decision-making, focusing on enabling humans to override mistaken AI advice [62, 29, 63, 25]. However, the proposed metrics are not readily applicable in the face of label indeterminacy, where the correctness of predictions is unknowable. In light of this, we suggest that our evaluation approach can serve as a blueprint for future studies: $(i)$ split the data into the set of instances with known labels and the label indeterminacy set; $(ii)$ evaluate models through traditional performance metrics, such as accuracy or AUC, on the set of instances with known labels (if any); $(iii)$ compare differences in predictions on the label indeterminacy set. An important open question for future work is how to define notions of human-AI complementarity and appropriate reliance in the presence of label indeterminacy.
@@
Reporting practices for AI-assisted decision-making typically contain an account of “what the AI system is capable of doing” and “how often the AI system may make mistakes” [1]. Making indeterminacy more explicit in reporting would significantly improve transparency in the deployment of AI systems for decision support by offering a clearer picture of their limitations and variabilities. Additionally, when true labels are not readily available and are instead constructed, it is crucial to report the choices and assumptions underlying this process. In particular, we recommend reporting $(i)$ methods used as part of the label construction, such as different forms of label aggregation or counterfactual estimation; $(ii)$ motivation or rationale underlying the method choice; $(iii)$ assumptions the method makes, and whether these are verifiable and indeed verified; and $(iv)$ arbitrary choices made throughout the construction process, such as uninterpretable parameters, which may impact the resulting labels and, thus, predictions. These dimensions extend existing reporting frameworks such as data cards [58], model cards [48], or datasheets for datasets [26]. When multiple methods for label construction are considered, these should all be reported, together with their implications for the resulting tools and corresponding performances.
@@
When designing AI systems for decision support, it is crucial to reconsider both the nature of AI predictions and their usage. Traditionally, AI systems focus on predicting specific outcomes, but our work shows how unreliable this approach is in the light of label indeterminacy. We encourage an increased focus on sociotechnical designs: rather than attempting to predict the outcome itself (e.g., recovery or loan default), one could design AI systems that predict a piece of information not suffering from indeterminacy and is useful—in combination with other complementary sources of information—to reach a high-quality decision [32]. For instance, AI systems could provide insights into how experts have evaluated similar cases in the past [19]. On these grounds, our work also has important implications for human-centered explainable AI [20] and research on cognitive forcing functions [8]. Rather than focusing solely on elucidating the “inner workings of AI,” research in these areas should prioritize mechanisms that incorporate social dimensions of decision-making [66]. Information concerning what construct is being predicted, and its differences with the construct of interest to the decision-making task, should be clearly communicated to users; otherwise, it could lead to a drift in decision-making criteria [28].
Moreover, incorporating uncertainty inherent to indeterminacy into predictions could enhance decision support; for instance, through informing deferrals or second opinions [42, 43, 49]. Instead of offering only a point prediction, systems might predict uncertainty based on whether experts have previously reached consensus on similar cases, or the degree to which individual predictions differ based on arbitrary design choices. Uncertainty is considered a key concept in decision support [40, 23, 76], and label indeterminacy provides a novel perspective for developing techniques to estimate and communicate this uncertainty.
6 Conclusion
@@ Appendix A Summary of All Models in Our Study
Table 4 summarizes the ten different approaches to estimating and incorporating unknown labels for WLST cases in our study. Recall that $\widetilde{Y}_{j}$ denote estimated labels, and $\varnothing$ refers to situations where labels are not augmented.
Table 4: Summary of all models and their respective label construction for WLST cases.
@@ Appendix B Additional Pairwise Comparison Tables
Table 5: Number of WLST patients with predicted likelihood to recover of $\leq$1in one model (rows) and $>$25% in another (columns).
@@
Table 6: Number of WLST patients with predicted likelihood to recover of $\leq$1in one model (rows) and $>$50% in another (columns).

### v2

@@ 4.5. Results: Predictions on Label Indeterminacy Set
Table 3. Number of patients with predicted likelihood to recover of $\leq$1in one model (rows) and $>$10% in another (columns). This shows that the ordering of predictions changes between models.
@@
Figure 5. Predictions for two WLST patients with strong disagreement between models. This disagreement is entirely due to unverifiable design choices made in estimating and incorporating labels for WLST patients.
[…]
Predictions of all ten models for two different patients. Model disagreement is very high.
Importantly, we find that the problem spans beyond calibration. In our study, it is not the case that predictions are merely “shifted” up or down from one model to another; the ordering of patients also changes. In other words, the difference between models is not only explained by different calibration. For instance, there are 35 patients for whom the model $M_{exp_{max}}$ predicts a recovery potential of $>$10% and $M_{obs}$ only $\leq$1%. These pairwise comparisons are shown in Table 3 for all models, as well as in Tables 5 and 6 in Appendix B for more extreme discrepancies. Overall, 323 WLST patients (approx. 19.6% of our sample) have a predicted likelihood to recover of $\leq$1by at least one model and $>$10% by at least one other. When comparing ranks, [^4] We look at the ranking of patients to study whether the differences in predictions across models can be explained as a calibration issue, which we confirm is not the case. We acknowledge that the decision of whether to withdraw life-sustaining therapies is typically not a ranking problem—unless one is operating under limited resources, such as allocating respirators during the COVID-19 pandemic (White and Lo, 2020). 74 patients are ranked in the bottom third by at least one model and in the top third by at least one other model in terms of their likelihood to recover. We visualize these stark differences in individual predictions in Figure 5 for two exemplary patients, where at least one model predicts a near-zero likelihood to recover and at least one other predicts $\gg$50%. In particular, the models providing optimistic vs. pessimistic predictions can vary significantly depending on the patient. For instance, $M_{exp_{max}}$ estimates a recovery chance of over 70% for patient PUH-2015-150 (Figure 5 left), while $M_{nn}$ suggests a recovery likelihood of less than 5% for the same patient. In contrast, when we examine patient PUH-2019-009 (Figure 5 right), the predictions are reversed. These differences—which have critical and potentially irreversible implications for patients—are entirely a result of unverifiable design choices made in handling indeterminate labels.
Crucially, when comparing the heterogeneity of model predictions between WLST and non-WLST (i.e., the ones where labels are known) patients, we infer from Figures 6 and 7 in Appendix C that models tend to disagree more for WLST patients—both in terms of prediction disparities and the number of patients receiving conflicting recovery prognoses. This shows that the disagreement we observe across models is not simply a function of predictive multiplicity. Moreover, the increased heterogeneity for WLST patients is especially problematic because the WLST cases are precisely the ones for which we do not have true labels and, as a result, cannot assess the correctness of predictions. This results in fundamental limitations for principled model selection.
5. Implications
@@
Deploying AI systems in high-stakes domains introduces specific ethical challenges that stem from label indeterminacy. A key concern is the arbitrariness of decisions these systems may inform, wherein different and unjustifiable methodological choices for handling indeterminate labels can yield outcomes lacking any principled grounding (Creel and Hellman, 2022; Black et al., 2022). For instance, a system might predict no chance of recovery for a patient under one label-handling method, while another method suggests a positive recovery potential for that same patient. In a healthcare context, this arbitrariness may influence whether life-sustaining therapies are withdrawn, potentially resulting in premature life-or-death decisions. This echoes concerns stemming from predictive multiplicity (Marx et al., 2020; D’Amour et al., 2022). While there is a similarity insofar as arbitrary choices yield different predictions, there is a differentiating factor: in the case of label indeterminacy, the issue does not stem from underspecification, under which one problem formulation with a known true label accepts multiple solutions (D’Amour et al., 2022). Instead, it stems from an ill-posed problem formulation, which relies on unknowable information, and in the face of it makes an arbitrary choice as to what that information may be. The “ground truth” is chosen to be one among many options, without knowledge of which one is more likely to be correct, but with full knowledge that at most one of them may be correct. As a result, the differences in predictions stem from some models relying on incorrect information. This poses unique ethical challenges for the design and deployment of these systems.
Crucially, in current practice there is often an implicit and unchallenged assumption that one method of estimating labels is correct, with alternative approaches rarely compared. This narrow focus obfuscates the indeterminate nature of the problem. Moreover, the opacity of assumptions implicit in this selection process can obscure the mechanisms behind AI recommendations. For patients and providers, this is particularly problematic as it veils an important dimension of uncertainty. This can, among other risks, foster an over-reliance on these recommendations (Buçinca et al., 2021; Schoeffer et al., 2025). Without a clear understanding of the uncertainties and limitations, patients and providers might trust the AI recommendations uncritically, potentially introducing new and consequential risks in care delivery. We refer to Mertens et al. (2021), among others, for additional ethical concerns related to the use of AI systems in neurological prognostication, which are beyond the focus of this work.
@@
This also carries implications for research and practice on human-AI collaboration (Fok and Weld, 2024; Bansal et al., 2021; Schemmer et al., 2023). Recent work in this field has advocated for a focus on appropriate reliance when evaluating AI-assisted decision-making, focusing on enabling humans to override mistaken AI advice (Schemmer et al., 2023; He et al., 2023; Schoeffer et al., 2024; Fok and Weld, 2024). However, the proposed metrics are not readily applicable in the face of label indeterminacy, where the correctness of predictions is unknowable. In light of this, we suggest that our evaluation approach can serve as a blueprint for future studies: $(i)$ split the data into the set of instances with known labels and the label indeterminacy set; $(ii)$ evaluate models through traditional performance metrics, such as accuracy or AUC, on the set of instances with known labels (if any); $(iii)$ compare differences in predictions on the label indeterminacy set. An important open question for future work is how to define notions of appropriate reliance in the presence of label indeterminacy.
@@
Reporting practices for AI-assisted decision-making typically contain an account of “what the AI system is capable of doing” (Amershi et al., 2019). Making indeterminacy explicit in reporting would significantly improve transparency in the deployment of AI systems for decision support by offering a clearer picture of their limitations. Additionally, when true labels are not readily available and are instead constructed, it is crucial to report the choices and assumptions underlying this process. In particular, we recommend reporting $(i)$ methods used as part of the label construction, such as different forms of label aggregation or counterfactual estimation; $(ii)$ motivation or rationale underlying the method choice; $(iii)$ assumptions the method makes, and whether these are verifiable and indeed verified; and $(iv)$ arbitrary choices made throughout the construction process, such as uninterpretable parameters, which may impact the resulting labels and, thus, predictions. These dimensions extend existing reporting frameworks such as data cards (Pushkarna et al., 2022), model cards (Mitchell et al., 2019), or datasheets for datasets (Gebru et al., 2021). When multiple methods for label construction are considered, these should all be reported, together with their implications for the resulting tools and corresponding performances.
@@
When designing AI systems for decision support, it is crucial to reconsider both the nature of AI predictions and their usage. Traditionally, AI systems focus on predicting specific outcomes, but our work shows how unreliable this approach is in the light of label indeterminacy. We encourage an increased focus on sociotechnical designs: rather than attempting to predict the outcome itself (e.g., recovery or loan default), one could design AI systems that predict a piece of information not suffering from indeterminacy and that is useful—in combination with other complementary sources of information—to reach a high-quality decision (Holstein et al., 2023). For instance, AI systems could provide insights into how experts have evaluated similar cases in the past (Ehsan et al., 2021). On these grounds, our work also has important implications for human-centered explainable AI (Ehsan et al., 2022) and research on cognitive forcing functions (Buçinca et al., 2021). Rather than focusing solely on elucidating the “inner workings of AI,” research in these areas should prioritize mechanisms that incorporate social dimensions of decision-making (Smart and Kasirzadeh, 2024). Information concerning what construct is being predicted, and its differences with the construct of interest to the task, should be clearly communicated to users; otherwise, it could lead to a drift in decision-making criteria (Green and Chen, 2019).
Moreover, incorporating uncertainty inherent to indeterminacy into predictions could potentially enhance decision support; for instance, through informing deferrals or second opinions (Lu et al., 2024; Madras et al., 2018; Mozannar et al., 2023). Instead of offering only a point prediction, systems might predict uncertainty based on whether experts have previously reached consensus on similar cases, or the degree to which individual predictions differ based on arbitrary design choices. Uncertainty is considered a key concept in decision support (Leibig et al., 2017; Fernandes et al., 2018; Zhang et al., 2024), and label indeterminacy provides a novel perspective for developing techniques to estimate and communicate this uncertainty.
6. Conclusion
@@
Acknowledgements.
This research was supported by NIH/NHLBI grant R01NS124642, by a Google Award for Inclusion Research, and by Good Systems, [^5] http://goodsystems.utexas.edu/ a UT Austin Grand Challenge to develop responsible AI technologies. The authors acknowledge the efforts of the Optimizing Recovery Prediction After Cardiac Arrest (ORCA) Study Group participants, who reviewed each case when death occurred after limitation or withdrawal of life-sustaining therapies to provide estimates of recovery potential had life-sustaining treatments been continued: Dr. Alain Cariou, Dr. Alejandro A. Rabinstein, Dr. Alexis Steinberg, Dr. Andrea O. Rossetti, Dr. Ankur A. Doshi, Dr. Bradley J. Molyneaux, Dr. Cameron Dezfulian, Dr. Carolina B. Maciel, Dr. Cecelia Ratay, Dr. Christoph Leithner, Dr. Cindy Hsu, Dr. Claudio Sandroni, Dr. Clifton W. Callaway, Dr. David M. Greer, Dr. David B. Seder, Dr. Francis X. Guyette, Dr. Fabio Silvio Taccone, Dr. Hiromichi Naito, Dr. Jasmeet Soar, Dr. Jean-Baptiste Lascarrou, Dr. Jerry P. Nolan, Dr. Jonathan Elmer, Dr. Karen G. Hirsch, Dr. Katherine Berg, Dr. Marion Moseby-Knappe, Dr. Markus B. Skrifvars, Dr. Michael Donnino, Dr. Michael Kurz, Dr. Min Jung Kathy Chae, Dr. Mypinder Sekhon, Dr. Nicholas J. Johnson, Mr. Patrick J. Coppler, Dr. Pedro Kurtz, Dr. Romergryko G. Geocadin, Dr. Sachin Agarwal, Dr. Teresa L. May, and Dr. Theresa Mariero Olasveengen.
@@ Appendix A Summary of All Models in Our Study
Table 4 summarizes the ten different approaches to estimating and incorporating unknown labels for WLST cases in our study. Recall that $\widetilde{Y}_{j}$ denotes estimated labels, and $\varnothing$ refers to situations where labels are not augmented.
Table 4. Summary of all models and their respective label construction for WLST cases.
@@ Appendix B Additional Pairwise Comparison Tables
Table 5. Number of WLST patients with predicted likelihood to recover of $\leq$1by one model (rows) and $>$25% by another (columns).
@@
Table 6. Number of WLST patients with predicted likelihood to recover of $\leq$1by one model (rows) and $>$50% by another (columns).
@@
Appendix C Prediction Heterogeneity for WLST vs. Non-WLST Patients
Figure 6 shows the distribution of the differences between the most optimistic and the most pessimistic recovery prognosis of any of the ten models $M_{obs},...,M_{exp_{agr_{w}}+ip}$ for a given patient (left) as well as the variance of different models’ predictions (right) for patients with known labels (Non-WLST) and those with unknown labels (WLST). We see that both the maximum differences and variance in predictions are larger for WLST patients. Figure 7 shows the percentage of patients for whom at least one model predicts a likelihood of recovery of $\leq$1and at least one other model predicts $>$10% (left), $>$25% (middle), or $>$50% (right), indicating stark discrepancies. Again, these discrepancies are significantly more pronounced for WLST patients. Taken together, this means that the heterogeneity resulting from different modeling approaches is particularly strong for WLST patients. This is especially problematic because those are precisely the instances for which we do not have true labels and, as a result, cannot assess the correctness of predictions.
Figure 6. Strip and box plots of maximum difference (left) and variance (right) between different models’ predictions for non-WLST (blue) and WLST (orange) patients.
[…]
Plots of maximum difference and variance between model predictions for non-WLST and WLST patients. Both metrics are higher for WLST patients.
Figure 7. Percentage of non-WLST (blue) and WLST (orange) patients for whom at least one model predicts a likelihood of recovery of $\leq$1and at least one other of $>$10% (left), $>$25% (middle), or $>$50% (right).
[…]
Percentage of non-WLST and WLST patients with significant discrepancies between different models’ predictions. Discrepancies are higher for WLST patients.